OpenStudy (abbycross167):

Can someone please check my work on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Solve the quadratic equation by factoring -- 3x^3 – 12x = 0

OpenStudy (abbycross167):

@jim_thompson5910 can you please help me?

OpenStudy (mathmale):

What quantity (represented by the product of a number and a letter) represents a common factor of the given quadratic equation? If you can identify that, then please write out that common factor as the multiplier of whatever is left. correct format: wx(y - z) = 0

OpenStudy (mathmale):

Be 100% certain that you know what "common factor" means in this context.

OpenStudy (abbycross167):

so would the common factor be x?

OpenStudy (abbycross167):

@agent0smith could you please try to help me?

OpenStudy (agent0smith):

Look at the quadratic. What common factor can you take out of both terms?

OpenStudy (abbycross167):

3x?

OpenStudy (agent0smith):

Yes.

OpenStudy (abbycross167):

okay then what do I do?

OpenStudy (agent0smith):

Take out the common factor out of both terms

OpenStudy (abbycross167):

what do you mean? like 3x^3 becomes 3x and 12x becomes 3x?

OpenStudy (agent0smith):

No. If you were factoring 5x^4 + 35x you'd write it as 5x(x^3 + 7)

OpenStudy (abbycross167):

I'm not sure how I would write 3x^3 – 12x = 0

OpenStudy (abbycross167):

@agent0smith ?

OpenStudy (agent0smith):

Try it.

OpenStudy (abbycross167):

3x(x^3 – 12) like that?

OpenStudy (agent0smith):

You took an x out of 3x^3 though... and a 3 out of 12. Try again

OpenStudy (abbycross167):

what do you mean?

OpenStudy (abbycross167):

3(x^3 – 9) thats what I needed?

OpenStudy (agent0smith):

No... taking out a factor is like dividing. You're doing it as if it's subtraction.

OpenStudy (abbycross167):

so 3x(x^2 – 3x)

OpenStudy (agent0smith):

Not that either...

OpenStudy (abbycross167):

okay, then idk what you mean...

OpenStudy (agent0smith):

You'll need to review some basics on factoring.

OpenStudy (abbycross167):

okay, but I was trying to write it in the format that you did this one 5x(x^3 + 7)